### v1

# A topological counting rule for shells

Abstract
Holding a shell in their hands, one can apply six loads: three by pulling and shearing, and three by bending and twisting. Here, it is shown that the shell will resist exactly three load cases and comply with the other three, provided the shell is simply connected, meaning it has no holes and no handles.
Maxwell [1] famously derived a counting rule for the zero modes of a truss. These are motions of the truss that do not stretch any of its members. The rule assigns one zero mode for each degree of freedom in excess of the number of drawn elastic bonds. Calladine [2, 3] refined the rule to account for redundant bonds: adding a bond either reduces indeterminacy by eliminating a zero mode or increases redundancy by creating a state of self-stress. The latter is a field of bar tensions in equilibrium with no external forces or reactions. Equivalently, a state of self-stress represents a way to redistribute tension from one bar to others, hence a redundancy.
For isostatic trusses, meaning with as many degrees of freedom as bonds, Kane and Lubensky [4] proposed a local rule that counts the zero modes supported by a subsystem. Unlike Maxwell–Calladine, theirs is sensitive to the mode shape, particularly to where and how the mode localizes along the subsystem’s boundary. Such sensitivity can only arise in sufficiently large trusses, where localization effects can manifest. Their rule also explains the emergence of robust, localized zero modes, analogous to protected electronic boundary states, using the topology of the spectrum of the compatibility matrix rather than the topology of the truss itself. Consequently, the Kane–Lubensky rule has powerful implications for the behavior of free surfaces in isostatic lattice materials, such as during indentation [5], crack initiation, or crack growth [6].
In cases where the truss triangulates a shell, its zero modes are discrete counterparts of the shell’s isometric deformations. These deformations, a classical subject in surface geometry, have regained prominence with advances in space deployable structures [7], soft robotics [8], 4D printing [9] and soft matter physics more broadly [10, 11], where they are used to geometrically encode or model morphing paths and buckled shapes. Conversely, the absence of such deformations is indicative of a structurally efficient shell, one that carries loads by surface tension, needless of bending stresses [12]. Beyond what is directly implied by Maxwell-Calladine, results reminiscent of counting rules for surfaces are rare. Notable ones include ($i$) rigidity results (e.g., for convex polyhedra [13]), where the count is zero; ($ii$) Cohn-Vossen’s counting argument for surfaces of revolution [14]; (iiiiii) the uniqueness of the mechanism of a nondegenerate mesh of planar quads [15]; and ($iv$) Tachi’s conjecture on the existence of a two-degree-of-freedom family of periodic mechanisms in generic periodic triangular meshes [16]. Other results provide local counts, e.g., for a single vertex or a neighborhood [17, 18], or derive isometric deformations for specific surfaces, e.g., origami tessellations and surfaces of translation [19, 20, 21, 22, 23].
Here, a new counting rule is established for simply connected shells and trusses that discretize them. These trusses are locally isostatic but topologically trivial in the Kane–Lubensky sense. Nonetheless, the simply connected topology enables a count of macroscopic isometric deformations, i.e., membrane-like or flexure-like deformations that are uniform up to a certain correction. The count is simple: there exist exactly three such isometric deformations. Conversely, a shell resists exactly three macroscopic loads, tension and bending combined, provided it is simply connected, irrespective of its composition or geometry. Note that for modes to be meaningfully classified as membrane, flexure, or macroscopic, the truss must be periodic or statistically homogeneous. If not, a version of the rule holds still but is not as amenable to interpretation. The proof relies on two familiar results, seldom used in tandem: the static-geometric analogy, and the Hill-Mandel lemma. Both are introduced next, followed by the proof and by a discussion of the role of topology.
Figure 1: The static-geometric analogy for macroscopic fields. Wide arrows are admissible stresses; thin arrows are isometric deformations: (a) in a plane, admissible longitudinal tension corresponds to isometric lateral bending; (b) in a singly corrugated membrane, an admissible longitudinal moment corresponds to isometric lateral extension; (c) in a multiply connected grid of tubes, admissible longitudinal tension has no corresponding isometric lateral bending.
Consider a shell with a midsurface $z=z(x_{1},x_{2})$, of Hessian $\mathbf{H}_{z}$, and with a statically admissible distribution of membrane stresses $\bm{\sigma}$, namely such that
$\operatorname{div}\bm{\sigma}=0,\quad\mathbf{H}_{z}:\bm{\sigma}=0.$ (1)
Letting $\bm{\sigma}=\operatorname{cof}\mathbf{H}_{\phi}$ derive from an Airy stress function $\phi$ eliminates the first equation and recasts the second into Pucher’s equation [24]
$\mathbf{H}_{z}:\operatorname{cof}\mathbf{H}_{\phi}=0.$ (2)
Now if a displacement $(\mathbf{u}\equiv(u_{1},u_{2}),w)$ is isometric, then the infinitesimal strain
$\bm{\varepsilon}\equiv\bm{\nabla}^{s}\mathbf{u}+\bm{\nabla}w\otimes^{s}\bm{\nabla}z$ (3)
vanishes meaning that $\bm{\nabla}w\otimes^{s}\bm{\nabla}z$ is a compatible planar strain. The compatibility equation, expanded and reduced, turns out to be identical to (2) with $w$ replacing $\phi$. This is the essence of the static-geometric analogy: a deflection is isometric if, as a stress function, it is statically admissible, and conversely. The analogy is well-known [25]; here, the novelty shall be in completing it with integral formulas for average strains and stresses. In particular, for a statistically homogeneous, say periodic, shell let
$\begin{bmatrix}u_{\alpha}\\ w\end{bmatrix}=\begin{bmatrix}E_{\alpha\beta}x_{\beta}+z\chi_{\alpha\beta}x_{\beta}\\ -\frac{1}{2}\chi_{\mu\nu}x_{\mu}x_{\nu}\end{bmatrix}+\text{periodic corrections}$ (4)
be the displacement field due to a macroscopic membrane strain $\mathbf{E}$ and a macroscopic bending strain $\bm{\chi}$. If the deformation is isometric, then
$\mathbf{E}_{o}=\left\langle{z\mathbf{H}_{w}}\right\rangle,\quad\bm{\chi}_{o}=-\left\langle{\mathbf{H}_{w}}\right\rangle,$ (5)
where $\left\langle{}\right\rangle$ averages over the 2D unit cell, or a representative surface element in the stochastic case. The subscript ‘$o$’ will be used to denote isometric strains and will be dropped when $\mathbf{E}$ and $\bm{\chi}$ are meant to be general strains. Similarly, for a statically admissible periodic membrane stress $\bm{\sigma}$ of Airy stress $\phi$, the macroscopic membrane and bending stresses are
$\bm{\Sigma}=\left\langle{\bm{\sigma}}\right\rangle=\left\langle{\operatorname{cof}\mathbf{H}_{\phi}}\right\rangle,\quad\mathbf{M}=\left\langle{z\bm{\sigma}}\right\rangle=\left\langle{z\operatorname{cof}\mathbf{H}_{\phi}}\right\rangle.$ (6)
Thus, the static-geometric analogy further matches $\operatorname{cof}\bm{\chi}$ and $\bm{\Sigma}$ as well as $\operatorname{cof}\mathbf{E}_{o}$ and $\mathbf{M}_{o}$: overall, it provides an isomorphism
$\begin{bmatrix}\bm{\Sigma}\\ \mathbf{M}\end{bmatrix}=\begin{bmatrix}\mathbf{0}&-\operatorname{cof}\\ \operatorname{cof}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{E}_{o}\\ \bm{\chi}_{o}\end{bmatrix}$ (7)
between the spaces of macroscopic isometric deformations and of macroscopic admissible stresses, where the $\operatorname{cof}$ operator plays the role of a $90^{\circ}$ rotation; see Figure 1.
The other main ingredient is the Hill-Mandel lemma [26]. This is a useful result of homogenization theory that describes some boundary conditions such that the virtual stress-strain work is the same if computed using the local fields or the macroscopic fields. Namely, for any admissible periodic $\bm{\sigma}$ and any admissible $(\mathbf{u},w)$ in the sense of (4), one has
$\left\langle{\bm{\sigma}:\bm{\varepsilon}}\right\rangle=\bm{\Sigma}:\mathbf{E}+\mathbf{M}:\bm{\chi}.$ (8)
The proof is straightforward and employs the divergence theorem in conjunction with periodicity (4), admissibility (1) and the compatibility conditions (3).
What is lacking here, to apply arguments à la Maxwell-Calladine, is to identify the spaces of macroscopic admissible stresses and isometric strains with the image and kernel of some structural matrix. In the present context, the appropriate matrix turns out to be the matrix of effective elasticity moduli $\mathbf{A}$. Define it by the minimization procedure
$\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}\equiv\min_{(\mathbf{u},w)}\left\langle{\bm{\varepsilon}:\mathbf{c}:\bm{\varepsilon}}\right\rangle$ (9)
under the constraints of equation (4), where $\mathbf{c}$ is the shell’s membrane stiffness tensor. The goal is to show that the spaces $\text{Im}\mathbf{A}$ and $\text{Ker}\mathbf{A}$ are precisely the spaces of admissible macroscopic stresses and isometric macroscopic strains.
It is rather clear that a macroscopic strain in the kernel of $\mathbf{A}$ is isometric for it has zero energy and therefore zero membrane strain. For the converse, the polarization identity of the bilinear form $\mathbf{A}$ is used: a macroscopic isometric strain $(\mathbf{E}_{o},\bm{\chi}_{o})$ can be superposed to an arbitrary strain $(\mathbf{E},\bm{\chi})$ without changing the strain energy, since the local membrane strain does not change either. By taking the difference, it comes that
$\begin{bmatrix}\mathbf{E}\\ \bm{\chi}\end{bmatrix}^{T}\mathbf{A}\begin{bmatrix}\mathbf{E}_{o}\\ \bm{\chi}_{o}\end{bmatrix}=0,\text{ for all }(\mathbf{E},\bm{\chi}),$ (10)
meaning that macroscopic isometric strains are in the kernel of $\mathbf{A}$. As for stresses, let $(\bm{\Sigma},\mathbf{M})$ be in the image of $\mathbf{A}$ produced by some $(\mathbf{E},\bm{\chi})$. Let $\bm{\sigma}=\mathbf{c}:\bm{\varepsilon}$ be the membrane stresses obtained by the minimization of strain energy (9). The stationarity conditions readily imply that $\bm{\sigma}$ is statically admissible and the Hill-Mandel lemma ensures that $\bm{\Sigma}=\left\langle{\bm{\sigma}}\right\rangle$ and $\mathbf{M}=\left\langle{z\bm{\sigma}}\right\rangle$. Conversely, let $(\bm{\Sigma},\mathbf{M})$ be an admissible macroscopic stress obtained by averaging an admissible periodic membrane stress $\bm{\sigma}$. Testing $\bm{\sigma}$ against isometric strains shows, by the Hill-Mandel lemma, that $(\bm{\Sigma},\mathbf{M})$ is in the orthogonal to the kernel of $\mathbf{A}$ and therefore, by symmetry of $\mathbf{A}$, in the image of $\mathbf{A}$.
In conclusion, $\text{Im}\mathbf{A}\oplus\text{Ker}\mathbf{A}$ provides a decomposition of the space of pairs of symmetric second order tensors into a direct sum of the spaces of admissible macroscopic stresses and isometric macroscopic strains, while the analogy previously showed that these spaces are isomorphic:
$\begin{gathered}\{(\mathbf{\Sigma},\mathbf{M})\}\oplus\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=S^{2}\times S^{2},\\ \{(\mathbf{\Sigma},\mathbf{M})\}\cong\{(\mathbf{E}_{o},\bm{\chi}_{o})\},\end{gathered}$ (11)
where $S^{2}$ is the 3D space of symmetric second order tensors. Taking the dimensions yields at last the counting rule
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}=3.$ (12)
The topology of the shell plays a crucial role that remained implicit in the proof presented above and that should be highlighted. On one hand, suppose the shell had handles, then there would be topological obstructions to integrating an admissible stress $\bm{\sigma}$ into an Airy stress function $\phi$. The static-geometric analogy, for averages, will then only encapsulate states of stress that derive from $\phi$ meaning that the rule needs to be weakened into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\leq 3\quad\text{(with handles)}.$ (13)
On the other hand, suppose the shell had holes, then there would be extra admissibility constraints for stresses in the form of zero tractions at boundaries. The analogy fails by leaving out isometries that do not reproduce this condition, that is isometries that cannot be extended over the holes or, equivalently, that deform the holes. The rule then weakens into
$\dim\{(\mathbf{E}_{o},\bm{\chi}_{o})\}\geq 3\quad\text{(with holes)}.$ (14)
Through the involved isomorphism, the rule endows the space of macroscopic isometric deformations with a natural symplectic structure: for any pair of modes $(\mathbf{E}_{i},\bm{\chi}_{i})_{i=1,2}$ in that space
$\mathbf{E}_{1}:\operatorname{cof}\bm{\chi}_{2}-\bm{\chi}_{1}:\operatorname{cof}\mathbf{E}_{2}=0.$ (15)
This is a restatement of the Hill-Mandel lemma (8) with the isomorphism (7) in mind. Such a relationship was recently derived using purely geometric tools [27, 28]. Here, the novelty is in the topological counting rule. The above identity reduces, when either mode is a pure membrane, e.g., $\bm{\chi}_{1}=\mathbf{0}$, or a pure flexure mode, e.g., $\mathbf{E}_{1}=\mathbf{0}$, to a remarkable identity between the in- and out-of-plane effective Poisson’s coefficient of the shell [29, 30, 31, 32, 33, 34]. Thanks to the construction of the analogy-induced isomorphism, it is possible to draw a whole series of exact relationships among the effective moduli, all summarized into
$\mathbf{A}\begin{bmatrix}\mathbf{0}&-\operatorname{cof}\\ \operatorname{cof}&\mathbf{0}\end{bmatrix}\mathbf{A}=\mathbf{0}.$ (16)
This relationship is more robust than the counting rule strictly speaking. Indeed, one can reasonably assume that the effective tensor $\mathbf{A}$ depends continuously on the local membrane stiffness $\mathbf{c}$ which can then be taken to 0 at places so as to create holes. By contrast, the above relationship can certainly fail in the presence of handles, in which case $\mathbf{A}$ can become full rank (e.g., as in a foldcore sandwich panels [35]).
Equation (16) is a “microstructure-independent” exact relationship. It is reminiscent of earlier works in the theory of composites [36]. More so, several such works employ what is referred to as a “translation”, i.e., a transformation that maps fields on the right side of a constitutive equation (e.g., $\bm{\varepsilon}$ in $\bm{\sigma}=\mathbf{c}:\bm{\varepsilon}$) onto fields on the left side of the same equation (e.g., $\bm{\sigma}$). In 2D elasticity, a “translation” is provided by none other than the $\operatorname{cof}$ operator, i.e., a pointwise $90^{\circ}$ rotation [37]. The same “translation” is also familiar to practitioners of the art of graphic statics [38, 39, 16]. Therein, the “translation” rotates admissible fields of bar tensions into compatible fields of infinitesimal rotations.
Examples are abundant where a shell is flexible out of plane and stiff in the plane, including trivially the plane, but also pyramidal trusses [40] and the Ron Resch origami pattern [41]. There are also many examples of shells with a unique pure membrane mode (i.e., with $\bm{\chi}_{o}=\mathbf{0}$) and two flexure modes including the Miura ori, the eggbox pattern and any singly corrugated shell [20]. Beyond that, there are no reported examples of shells with two or three pure membrane modes. It is not known if there is a fundamental limit on the number of pure membrane modes, but this appears to be unlikely. Here, the unit cell of a shell with approximately three pure membrane modes is produced by numerical minimization of the trace of the effective membrane stiffness. While not a proof, the numerical experiment convincingly shows that it is possible for a periodic simply connected membrane to be stiff out of plane. The simulations employ a truss code with periodic boundary conditions to compute the effective membrane stiffness along with a standard minimization routine to minimize its trace. See  [^1] Code for membrane stiffness minimization is available at https://github.com/nassarh/ for implementation details; see also [43]. Figure 2 shows the results. Therein, it is remarkable that the optimized geometries achieve a reduction of the effective membrane stiffness by 10 orders of magnitude, with as little as $20\%$ changes in elevations, and no noticeable qualitative changes in the geometry. This reinforces the point of view that, however insightful, the present methods leave important quantitative questions unanswered, e.g., regarding the sensitivity of the space of isometries to perturbations.
Figure 2: Minimization of the effective membrane stiffness for discrete shells with (a) $2\times 2$, (b) $4\times 4$, and (c) $6\times 6$ nodes per unit cell. Random initial geometries are in the top row; geometries with optimized elevations are in the bottom row. The trace of the effective membrane stiffness tensor is reduced by a factor of $\sim 2$, of $\sim 10^{10}$ and of $\sim 10^{10}$, respectively.
The results presented here establish a macroscopic counting rule for macroscopic isometric deformations of simply connected shells. By combining the static–geometric analogy with the Hill–Mandel lemma, the spaces of macroscopic isometries and of admissible macroscopic stresses were proven to be complementary and isomorphic. Their dimensionality then yields precisely three macroscopic isometric modes, just as it yields precisely three admissible macroscopic stresses. The same arguments reveal an exact algebraic relationship between the effective elasticity moduli providing a microstructure-independent constraint on the homogenized constitutive law. Note that the fact that the shell’s surface was taken to be a smooth graph is a matter of convenience; The same techniques can be adapted to handle jumps in the tangent plane (e.g., straight and curved creases) as well as overhangs; see, e.g., the classical tools used in [28] as well as the measure theory formalism used in [44, 45]. By contrast, simple connectedness is an essential, both sufficient and necessary, ingredient of the theory. Indeed, topology is known to severely impact structural stiffness, as it does for instance in Saint-Venant’s problem of torsion for simply or multiply connected sections; the present context is no exception.
Acknowledgements.
Work supported by the NSF under CAREER award No. CMMI-2045881. Part of this work was inspired by a series of talks given at two symposia: Geometry of Materials at ICERM (Brown University) and GASM at IMAG (Universidad de Granada); the author thanks the speakers and organizers.